Circular No. 7902

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 K1 K. Lawrence, E. Helin, and S. Pravdo, Jet Propulsion Laboratory, re-

port the discovery of a comet with a tail ∼ 8′′ long toward the southwest (and m_{2} = 19.4–19.6) on images taken with the 1.2-m NEAT telescope at Haleakala. Confirming CCD images by E. S. Barker at McDonald Observatory (0.76-m reflector) show a 15′′ tail toward the southwest.

2002 UT α_{2000} δ_{2000} m_{1} Observer
May 16.54720 19^{h}56^{m}30.^{s}38 −24^{o}26′34′′.9  NEAT
16.55770 19 56 29.75 −24 26 26.1  _{”}
16.56867 19 56 29.08 −24 26 17.3  _{”}
17.42035 19 55 38.16 −24 14 15.8 17.9 Barker
17.42879 19 55 37.56 −24 14 09.2 18.1  _{”}

SUPERNOVA 2002cw IN NGC 6700 M. Papenkova and W. D. Li, University of California at Berkeley, re-

port the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on an unfiltered CCD image taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on May 16.5 UT (mag ∼ 18.7). The new object was confirmed on an earlier image taken on May 4.5 at about the same brightness. SN 2002cw is located at α = 18^{h}46^{m}02^{s}.40, δ = +32^{o}17′03′′.2 (equinox 2000.0), which is 25′′.0 west and 16′′.0 north of the nucleus of NGC 6700. A KAIT image taken on 2001 Oct. 1.2 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2002cx IN ANONYMOUS GALAXY W. M. Wood-Vasey, G. Aldering, and P. Nugent, Lawrence Berkeley

National Laboratory, report the discovery of an apparent supernova (mag 18.9) on unfiltered CCD images taken with the Oschin 1.2-m telescope at Palomar on May 12.21 UT as part of a joint collaboration with E. F. Helin, S. Pravdo, M. Hicks, and K. Lawrence (Near-Earth Asteroid Tracking project). SN 2002cx is located at α = 13^{h}13^{m}49^{s}.72, δ = +6^{o}57′31′′.9 (equinox J2000.0), which is 18′′ south and 11′′ east of the center of its apparent host galaxy. The new object was confirmed (mag 18.1) on subsequent images taken on May 16.2. No object is apparent on images from Apr. 23 and May 6 (limiting mag 20).

2002 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green